### v1

# Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration

Abstract
The evolution of artificial intelligence (AI) and neural network theories has revolutionized the way software is programmed, shifting from a hard-coded series of codes to a vast neural network. However, this transition in engineering software has faced challenges such as data scarcity, multi-modality of data, low model accuracy, and slow inference. Here, we propose a new network based on interpolation theories and tensor decomposition, the interpolating neural network (INN). Instead of interpolating training data, a common notion in computer science, INN interpolates interpolation points in the physical space whose coordinates and values are trainable. It can also extrapolate if the interpolation points reside outside of the range of training data and the interpolation functions have a larger support domain. INN features orders of magnitude fewer trainable parameters, faster training, a smaller memory footprint, and higher model accuracy compared to feed-forward neural networks (FFNN) or physics-informed neural networks (PINN). INN is poised to usher in Engineering Software 2.0, a unified neural network that spans various domains of space, time, parameters, and initial/boundary conditions. This has previously been computationally prohibitive due to the exponentially growing number of trainable parameters, easily exceeding the parameter size of ChatGPT, which is over 1 trillion. INN addresses this challenge by leveraging tensor decomposition and tensor product, with adaptable network architecture.
1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 has facilitated significant advancements in the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies within the field of engineering and science introduces distinct challenges. These challenges include scarcity and multi-modal aspect of data, low generalization error compared to classical numerical methods [4, 5, 6], slow inference (function evaluation) time that challenges real-time control [7, 8, 9, 10] and statistical modeling, and resource-heavy model architecture [11, 12, 13, 14].
In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of computational and data-driven engineering software into a new paradigm.
“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibration in science/engineering problems."
As one promising enabler of Engineering Software 2.0, we introduce interpolating neural networks (INN). The key ideas behind INN are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation points, 2) interpolate them with well-established interpolation techniques, and 3) optimize the values and coordinates of the interpolation points. This is distinct from the way it is referred to in the machine learning (ML) community; interpolation in ML generally refers to the interpolation of training data, which may cause overfitting. Figure 1b-d clearly shows the uniqueness and simplicity of our approach: interpolation of points (red dots in Figure 1d). Note that one can also extrapolate for a certain range if the interpolation points are located outside of the range of training data.
INN resolves the exponentially growing computational cost of high-dimensional interpolation by leveraging the separation of variables and tensor decomposition [15, 16, 17, 18] (see Figure 1e). INN facilitates three major tasks in computational science and engineering: 1) train (or learn) on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solve extremely challenging high-dimensional problems with statistics involved, and 3) calibrate (or solve inverse problems) for uncertain parameters.
Our results demonstrate that INN can markedly enhance training/inference speed and model accuracy compared to the feed-forward neural network (FFNN) or physics-informed neural network (PINN)[19] by dramatically reducing the number of trainable parameters and network connectivity. This also results in a huge reduction in the usage of random-access memory (RAM) and disk storage, facilitating the transferability of a trained/solved model across multiple workstations. Furthermore, INN can adapt its network structure to further improve model accuracy or to incorporate statistical behavior. INN has a huge potential of replacing FFNN and convolutional neural network (CNN) [20] in modern AI models with the aforementioned benefits.
Figure 1: Overview of Interpolating Neural Networks (INN). a, INN forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibration. The trainer employs data-driven loss functions (e.g., mean squared (MSE) error loss) while the solver adopts a residual loss of a partial differential equation (PDE). A trained/solved INN model can then be employed as a forward model of a calibrator to solve an inverse problem. It can also predict output uncertainties originating from input variables and the model itself. b-c, One-dimensional (1-D) input INN discretized with 5 segments and 6 interpolation points, where c is an expansion of segment 2. The purple lines denote a logical operation that sends an input variable $p$ to only one segmental network (black solid boxes) that it belongs to. The first two layers (green dotted boxes) build linear interpolants and the third layer (red dotted box) builds nonlinear interpolants ($\tilde{\bm{N}}(p)$). Finally, the outputs of the last hidden layer are dot-produced with corresponding point values $(\bm{u})$ to yield the interpolated output, $\bm{u}^{h}(p)$. d, a regression problem [21] trained with b-c, highlighted for segment 2. e, INN for multi-dimensional ($I$-D) inputs. Each network ($m=1,\cdots,M$) has $I$ different 1-D INNs illustrated in b, followed by a product operator $\prod_{i=1}^{I}$. The outputs of all networks are added to yield the interpolated output, $\bm{u}^{h}(\bm{p})$.
2 INN: Unifying Training/Learning, Solving, and Calibration
Interpolating Neural Network (INN) is a systematically pruned neural network based on interpolation theories [17, 22] that leverage classical numerical methods such as finite element [23] and meshfree methods [24, 25]. Although these numerical methods and the neural network theory have been developed by seemingly different groups of researchers, they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. INN is built upon this similarity and it forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibration, as illustrated in Figure 1a.
2.1 INN for Training
The key idea of INN is that the domain of an input variable is divided into non-overlapping segments with interpolation points, interpolation functions (or interpolants, which will be used interchangeably) are built for each segment, then the interpolation points (both values and coordinates) are optimized. An illustrative regression problem is shown in Figure 1d. Assume that the one-dimensional input domain is uniformly divided into 5 segments with 6 interpolation points whose coordinates ($p_{1}\sim p_{6}$) are fixed for now. For a given input data ($\prescript{*}{k}{p}$), we first find out which segment the input data belongs to; the input data $\prescript{*}{k}{p}$ is in segment 2. Then a patch domain is considered that encompasses neighboring segments, where the size is determined by patch size, $s$. Since the patch domain of segment 2 consists of five interpolation points ($u_{1}\sim u_{5}$), five INN interpolants $(\tilde{N}_{j}(p))$ are constructed to nonlinearly interpolate at $p=\prescript{*}{k}{p}$ as:
$u^{h}(\prescript{*}{k}{p})=\begin{bmatrix}\tilde{N}_{1}(\prescript{*}{k}{p})&\tilde{N}_{2}(\prescript{*}{k}{p})&\tilde{N}_{3}(\prescript{*}{k}{p})&\tilde{N}_{4}(\prescript{*}{k}{p})&\tilde{N}_{5}(\prescript{*}{k}{p})\end{bmatrix}\cdot\begin{bmatrix}u_{1}\\ u_{2}\\ u_{3}\\ u_{4}\\ u_{5}\end{bmatrix}.$ (1)
See Supplementary Section for detailed construction of INN interpolants.
The corresponding neural network structures are illustrated in Figure 1b-c and Figure2a for segment 2. INN’s weights and biases are functions of coordinates ($p_{i}$) and values ($u_{i}$) of the interpolation points, and INN hyper-parameters (e.g., patch size $s$, dilation parameter $a$). Assuming that the coordinates and hyper-parameters are fixed (i.e. INN interpolants are fixed), the only trainable parameters are the interpolation point values ($u_{j}$), which appear as weights between the third hidden layer and the output layer (see Figure 1c). For this reason, INN can dramatically reduce trainable parameters compared to FFNN, yielding faster training, less memory requirements, and faster inference. It is worth noting that the neural network structure of INN is not limited to Figure 1b; one can build INN with the first two hidden layers only for linear interpolation, or use other interpolation techniques such as meshfree [24, 25] or splines [26] that may result in different structures.
The goal of training for a regression problem is to minimize the distance between the prediction $u^{h}(\prescript{*}{k}{p})$ and the output data $\prescript{*}{k}{u}$. Thus, a mean-squared error (MSE) loss function is used:
$loss(\bm{u})=loss(u_{1},u_{2},\cdots,u_{6})=\frac{1}{K}\sum_{k}(u^{h}(\prescript{*}{k}{p})-\prescript{*}{k}{u})^{2},$ (2)
where $K$ is the number of training data. Finally, an optimization problem is formulated as:
$\underset{\bm{u}}{\text{minimize}}\ loss(u_{1},u_{2},\cdots,u_{6}).$ (3)
Figure 2: a, A sub-neural network that constructs convolution patch functions (weights between the second and third hidden layers, $\mathcal{W}$). Weights and biases in this network are functions of coordinates of interpolation points ($p_{i}$) and INN hyper-parameters. See supplementary Information for detailed formulation. b-c, Effect of activation functions when a sinusoidal ground truth function is trained. Details can be found in Supplementary Information. c, A trained INN with sinusoidal activation function showing a perfect fit to the ground truth function. d-e, convolutional neural network (CNN) version of INN where each network corresponds to a CNN filter.
We can also update the interpolants (i.e., the coordinates of interpolation points and INN hyper-parameters). For example, one can optimize $p_{1}\sim p_{6}$ to optimize the positions of the interpolation points (or the length of each segment) by:
$\underset{\bm{u}}{\text{minimize}}\ loss(p_{1},\cdots,p_{6};u_{1},\cdots,u_{6}),$ (4)
but still, the training cost is significantly lower than FFNN. One can also update the number of interpolation points adaptively, as discussed in Supplementary Information.
Similar to the activation function that determines the nonlinearity of FFNN, the convolution patch functions, appearing as the weights between the second and third hidden layers $(\bm{\mathcal{W}})$ in Figure 1c, control nonlinear interpolation in INN. The formulation of convolution patch functions with arbitrary activation is schematically illustrated in Figure 2a and discussed in detail in Supplementary Section. As the activation function dictates the interpolation basis, there might be a better- or worse- performing activation function depending on the ground truth function to be trained. Two illustrative examples are provided in Figure 2c,e. In general, activation functions close to the ground truth function yield low errors. Especially in Figure 2d where the two are the same, INN can reach a perfect fit to the ground truth function. Thus, activation functions are tunable hyperparameters. One can adaptively optimize them for each segment during training, which is prohibitive or less common in most neural networks [27, 28] (see Supplementary Section for detailed demonstration).
The INN interpolants can also be formulated for multi-dimensional inputs. In references [17, 22], which are the predecessors of INN theory, the benefits of those interpolants for solving 2D and 3D partial differential equations (PDEs) are demonstrated. However, the computational cost of building multi-dimensional interpolants grows exponentially with the input dimension. To tackle this problem, we propose a projected interpolation to INN, mathematically inspired by the separation of variables, proper orthogonal decomposition [29], proper generalized decomposition [30], and tensor decomposition [15].
The projected interpolation converts $n$-dimensional interpolation into “$n$” 1-dimensional interpolations with “$M$” different modes. With it, one can dramatically (typically by orders of magnitude) reduce the degrees of freedom (DoFs) of high-dimensional problems. Consider an $I$-input and $L$-output system. Similar to the single-input case, we discretize each input dimension and find the values of interpolation points through optimization. The predicted value at $k$-th input data $\prescript{*}{k}{\bm{p}}$ using INN becomes:
$\bm{u}^{h}(\prescript{*}{k}{\bm{p}})$ $=\bm{u}^{h}(\prescript{*}{k}{p}_{1},\prescript{*}{k}{p}_{2},\cdots,\prescript{*}{k}{p}_{I})=\sum_{m=1}^{M}[\tilde{\bm{N}}_{p_{1}}(\prescript{*}{k}{p}_{1})\cdot\bm{u}_{p_{1}}^{(m)}]\odot[\tilde{\bm{N}}_{p_{2}}(\prescript{*}{k}{p}_{2})\cdot\bm{u}_{p_{2}}^{(m)}]\odot\cdots\odot[\tilde{\bm{N}}_{p_{I}}(\prescript{*}{k}{p}_{I})\cdot\bm{u}_{p_{I}}^{(m)}]$ (5)
$=\sum_{m=1}^{M}\prod_{i=1}^{I}[\tilde{\bm{N}}_{p_{i}}(\prescript{*}{k}{p}_{i})\cdot\bm{u}_{p_{i}}^{(m)}]$
where $\prescript{*}{k}{\bm{p}}=[\prescript{*}{k}{p}_{1},\prescript{*}{k}{p}_{2},\cdots,\prescript{*}{k}{p}_{I}]$ is the input variables of $k$-th training data, $\bm{u}^{h}$ is the corresponding interpolated output vector of length $L$, $M$ is the number of modes (or networks; we use them interchangeably in INN), $\tilde{\bm{N}}_{p_{i}}(\ast)$ and $\bm{u}_{p_{i}}^{(m)}$ are the one-dimensional INN interpolants and values of interpolation points for the input dimension $p_{i}$, respectively, and $\odot$ refers to the component-wise multiplication of the same dimension. Starting from Eq. 5, $p_{i}$ will denote $i$-th input variable or dimension, and $p_{i,j}$ will refer to the $j$-th coordinate in $i$-th input dimension. If all input dimensions are discretized with the same number of segments (or interpolation points, denoted as $J$), the values of mode $1$ through mode $M$ become a 4-th order tensor of size $(M,I,J,L)$, which is the trainable parameters of INN. See Supplementary Section for detailed training algorithms.
The neural network structure of the multi-dimensional INN is illustrated in Figure 1c. Each mode in Eq. 5 is represented by an independent neural network with three hidden layers. Each network has the potential to possess unique hyperparameters, including the number of segments and activation function, thereby creating a new realm of significant versatility. For instance, one can start with a high number of segments for earlier networks and gradually reduce the level of discretization for later networks, mimicking multi-resolution filtering in wavelet analysis or Fourier transform [31], which has not been realized in previous works [17, 18].
It is also possible to nullify some input variables mode by mode, which is similar to the concept of random dropout in neural networks [32] or the convolutional neural networks (CNN) [20]. The method is simple: set interpolation point values of an input variable $p_{i}$ as 1, $u_{p_{i}}^{(m)}=1$, like applying Dirichlet boundary conditions for solving partial differential equations. As the interpolation of network $(m)$ is a product of "$I$" one-dimensional interpolations, this will result in dropping out or nullifying these inputs. One special case of manual dropout resembling CNN is illustrated in Figure 3d-e. Given a 5 $\times$ 5 pixels image, instead of having the entire 25 pixels as input variables, one can construct each network such that it takes inputs of one CNN filter. When a 3 $\times$ 3 filter is used with a stride of 1, each INN network gets 9-pixel inputs and the total number of networks becomes $M=9$. We conducted benchmarks on the fashion-MNIST dataset [33] with this idea which will be discussed in the following section.
2.2 INN for solving
The INN solver uses the same neural network structure as the trainer but is equipped with a loss function formulated from an equation to be solved. Here, we introduce a general formulation for solving parameterized time-dependent partial differential equations (PDEs), which is prohibitive for most numerical methods and machine learning approaches [34]. We construct an optimization problem from the PDE, which has been introduced as the residual minimization strategy [35] or the Petrov-Galerkin loss function [36, 37].
The approximate solution field of INN for a space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ problem is given as:
$\begin{split}\bm{u}^{h}(\bm{x},t,\bm{\theta})&=\sum_{m=1}^{M}[\tilde{\bm{N}}_{x_{1}}(x_{1})\cdot\bm{u}_{x_{1}}^{(m)}]\odot\cdots\odot[\tilde{\bm{N}}_{x_{d}}(x_{d})\cdot\bm{u}_{x_{d}}^{(m)}]\odot[\tilde{\bm{N}}_{t}(t)\cdot\bm{u}_{t}^{(m)}]\\ &\odot[\tilde{\bm{N}}_{\theta_{1}}(\theta_{1})\cdot\bm{u}_{\theta_{1}}^{(m)}]\odot\cdots\odot[\tilde{\bm{N}}_{\theta_{k}}(\theta_{k})\cdot\bm{u}_{\theta_{k}}^{(m)}].\end{split}$ (6)
where $d$ and $k$ are the spatial dimension and the number of parameters, respectively. Similar to the trainer, the goal is to find values of interpolation points: $\bm{u}^{(1\sim M)}_{\bm{x}},\bm{u}^{(1\sim M)}_{t},\bm{u}^{(1\sim M)}_{\bm{\theta}}$. The loss function is derived from weighted integrals of the differential equation ($loss^{PDE}$) and boundary ($loss^{BC}$)/initial ($loss^{IC}$) conditions:
$loss$ $=loss^{PDE}+\tau_{BC}\cdot loss^{BC}+\tau_{IC}\cdot loss^{IC}$ (7)
See Supplementary Section for detailed derivations of the loss function, test function, and numerical integration. The weights $\tau_{BC}$ and $\tau_{IC}$ control the rate of convergence and can be user-specified.
2.3 INN for calibration
[…]
$\prescript{*}{}{\bm{p}}=\underset{\bm{p}}{\text{argmin}}\frac{1}{K}\sum_{k}(\bm{u}^{h}(\bm{p})-\prescript{*}{k}{\bm{u}})^{2},$ (8)
where $\bm{u}^{h}(\bm{p})$ is the trained or solved forward model, $\prescript{*}{k}{\bm{u}}$ is the $k$-th measured observation and $\prescript{*}{}{\bm{p}}$ is the calibrated input variables.
3 Benchmark Demonstration
To demonstrate the advantages of INN for training, calibration, and solving, we present several benchmarks for each subject. The benchmark problems include training of data that follow scientific equations, classification of spiral data, classification of the fashion-MNIST dataset [33], calibration of parameters in the scientific equations and electrode manufacturing data [38], and solving of a parameterized partial differential equation (PDE). Except for the fashion-MNIST classification, all benchmark demonstrations are provided in the supplementary information, clearly revealing significant benefits such as less trainable parameters, high accuracy, less memory requirement, fast training, and inference time. One can also find how the INN hyperparameters are selected in Supplementary Information.
Table 1 shows the benchmark result of the fashion-MNIST image classification problem [33]. Results of the support vector classifier (SVC), random forest classifier (RFC), and tensor neural network (TNN) are taken from references, while FFNN, INN, and INN (CNN type) are from our own source code. Compared to well-established classifiers (SVC, RFC, FFNN), INNs can achieve comparable high accuracy with fewer trainable parameters. Especially, the CNN-type INN can only use around 50% parameters of FFNN without significantly sacrificing the prediction accuracy. It is worth noting that the TNN, which shares a similar mathematical background with INN (i.e., tensor decomposition and contraction), is less accurate than all the others and limited to smaller image size [39].
Table 1: Benchmark result of the fashion-MNIST classification.
Support Vector Random Forest Tensor Neural FFNN INN INN
Classifier [33] Classifier [33] Network [39] (CNN type)
Features RBF kernel Estimators: 100 Reduced image [784,512,256,10] 30 networks 5$\times$5 filter
Gamma:10 Max depth: 100 8$\times$8 pixels neurons 1 segments stride: 1
Test 0.8970 0.8730 0.8343 0.8918 0.8813 0.8731
accuracy
4 Scientific and Engineering applications
Figure 3: Application of the INN methods to scientific and engineering problems. During solving a very high-dimensional manufacturing process of L-PBF, the INN method is $10^{5}$ and $10^{6}$ times faster than FEM and PINN. For heat source calibration, INN requires very high-fidelity data to develop an accurate surrogate model. To develop a data-driven real-time online monitoring tool, INN uses only 18% training parameters and is 18-31 times faster compared to FFNN.
This section will illustrate how Engineering Software 2.0 can be applied to model various facets of metal additive manufacturing by integrating training, solving, and calibration processes. Laser Powder-bed Fusion (L-PBF) is a type of metal additive manufacturing in which metal powders are layered and melted by a laser beam to create objects that match the provided CAD designs [40]. The added flexibility of L-PBF comes with very large design space and spatially varying microstructure after manufacturing. Therefore, considerable research is being devoted to employing in situ monitoring data for real-time control of L-PBF processes to minimize variability [7]. The modeling aspect of L-PBF spans across modeling the manufacturing process with very high design space (solving), computing the internal relationship among different parameters of manufacturing (calibrating), and designing an online monitoring and control system for L-PBF (training).
The governing equation for modeling the manufacturing processing is given by
$\frac{\partial T(\mathbf{x},t)}{\partial t}-k\Delta T(\mathbf{x},t)=Q_{h}(x,y,t,P,R,\eta,d)$ (9)
This is a heat conduction equation where the moving laser heat source is incorporated through the source term, $Q_{h}$. The equation involves spatial parameter $(x,y,z)$, temporal variable $t$, thermal conductivity $k$, power $P$, beam radius $R$, absorptivity $\eta$, and laser depth $d$. For the simplest case of single track laser pass, the initial condition is $T(\mathbf{x},0)=0$, and the boundary conditions are $T=\tilde{T}$ in $\partial\Omega_{T}$, and $\frac{\partial T}{\partial\mathbf{x}}\cdot\mathbf{n}=\bar{q}$ in $\partial\Omega_{q}$.
4.1 Solving High Dimensional Space-Time-Parameter Heat Equation
Equation 9 has 9 input variables; 3 for spatial domain, 1 for time domain, and 5 for parametric domain. See, Supplementary Section for a detailed problem definition. This equation is prohibitive for process design with most existing computational methods, including finite element method (FEM)-based data-driven method and physics informed neural network (PINN) [19].
To solve the problem with FEM-based data-driven method, one first needs to sample from the 5-dimensional parametric domain, followed by running space-time FEM for each sampled parameter set. Then the FEM solution should be stored as training data, and a neural network surrogate model will be trained based on the data. The computational cost required is prohibitive as illustrated in Supplementary Section. Moreover, as a larger number of training data are generated due to the high dimensionality of the input domain, this method becomes prohibitive in terms of data storage. Running large-scale FEM simulation is also memory inefficient.
PINN is a unified framework of training and solving such parameterized PDEs, however, the computational cost still grows exponentially as the number of input variables grows. As illustrated in Supplementary Section, it is almost impossible to solve this 9-dimensional equation. In addition, the convergence of solution accuracy is not guaranteed, making it skeptical if PINN can even find a good solution for more than 7 input variables. PINN also suffers from memory issues for GPU training due to the exponential growth of the collocation points (although batch training can be adopted to alleviate memory usage).
Contrary to FEM and PINN, INN solves this space-time-parameter problem all at once. As it is based on the projected interpolation, the trainable parameters linearly grow with the input variables, making the computational cost manageable. Meanwhile, since tensor-decomposed data structures are leveraged, INN has the benefits of significant memory efficiency and storage gain especially for high-dimensional problems. A detailed comparison is provided in Supplementary Section.
On another note, once the parametric solution of the AM problem is obtained, we can easily investigate uncertainty propagation by imposing probabilistic distributions on the inputs of INN, thanks to the tensor-decomposed data structure. Figure 3 shows the uncertainty of the output when the inputs are assumed as independent Gaussian variables (see details in Supplementary Section).
4.2 Calibration of Heat Source Parameters for Additive Manufacturing
In this example, the INN is used for calibrating a model of heat source of L-PBF. Usually, a Gaussian beam profile is assumed as a volumetric heat source in L-PBF modeling [41]. The heat source equation reads,
$Q_{h}(\textbf{x},t,\textbf{p})=\frac{2\mathcal{P}\eta}{\pi r^{2}_{b}d}e^{-\frac{2(x^{2}+y^{2})}{r^{2}_{b}}}$ (10)
where, $z_{top}-z\leq d$. The notations follow Eq. 9. To get an accurate solution, some parameters need to be calibrated against experimental data. For example, three calibration parameters $p_{1}$, $p_{2}$, and $p_{3}$ can be defined as, $d=p_{1}\frac{\mathcal{P}}{V}$, $\eta=p_{2}\frac{\mathcal{P}}{V}$, and $r_{b}=p_{3}\frac{\mathcal{P}}{V}$. Here, $V$ is the scan speed. However, the challenges for calibrating these parameters are: 1) the computational models are often very expensive to perform even one forward run, 2) the computational models are not automatic differentiable, and 3) limited calibration data is available. As a demonstration for this problem, INN is trained for 7 sets of meltpool temperature data generated by FEM. Later, this trained INN is used for optimizing the three aforementioned parameters using gradient-based optimizers. Even with only 7 sets of simulation data, the INN model could generate calibrated parameters that produce meltpool temperature (or TEP) within 6.4% of the experimental data (see Figure 3).
4.3 Real-time Online Control of Laser Powder Bed Fusion Additive Manufacturing
In this application, the interpolating neural network is used to devise an online control system for L-PBF additive manufacturing. The goal is to maintain a homogeneous melt pool temperature. The hypothesis is a homogeneous melt pool temperature across each layer of material deposition will reduce the variability in the microstructure and mechanical properties. The fundamental challenge of implementing a model predictive control system for such a system is the computational time required for the forward prediction (involving finite element or computational fluid dynamics models), and the lack of automatic differentiation capabilities in the available computational models. The contemporary data-driven models require a significant amount of data and training parameters for accurate prediction which in turn limit the applicability of these methods for real-time prediction. Therefore, a fast, memory-efficient, and differentiable computational method is required.
For prediction, the data is generated by fusing experiments and computational methods (see Supplementary Section for details). The online control loop uses the INN model to inform the manufacturing machine. There are two models: a) the forward model uses the mechanistic features (precomputed based on the tool path and geometry of the build to generalize a data-driven model for unseen tool path and geometric features; more details can be found in [42]), and melt pool temperature of the previous layer (measured by the sensor as Thermal Emission Plank (TEP) signal; more details in the supplementary information) to predict the TEP/melt pool temperature in the next layer for an uncontrolled process. The uncontrolled TEP is converted to a homogenized TEP over the entire layer. b) the inverse model takes this targeted, homogenized TEP as the input along with the previous layer’s mechanistic features to predict the required laser power for the next layer. The laser power is then exported to the machine to print the next layer. The forward model is trained on experimentally observed data while the inverse model is trained on computational data coming from a finite element solver. The INN model is applied for both forward and inverse prediction and compared against FFNN. Figure 3 shows that for the inverse model INN is at least 18 to 31 times faster to train compared to the FFNN to reach the same level of training error, and it uses only about 18% of the parameters that FFNN uses. The speed-up depends on the amount of data used. The INN method is consistently more accurate compared to FFNN (see Supplementary Section for more details).
5 Conclusions
We show that interpolating neural networks (INN) can train, solve, and calibrate scientific and engineering problems that are extremely challenging or prohibitive for existing numerical methods or machine learning models. The keys to INN’s success in those problems are 1) it interpolates virtual points with well-established interpolation theories (not interpolating training data with a black-box network) and 2) it leverages the idea of separation of variables and tensor decomposition to resolve the curse of dimensionality. Due to the significantly reduced trainable parameters without losing approximation capabilities, INN can train faster with less memory requirement, and achieve extremely high model accuracy and fast inference time.
As this is the original article on INN, we have myriads of research questions that may gather huge research interests across various areas including but not limited to scientific machine learning, applied mathematics, computer vision, and data science. To mention a few:
  • Multi-resolution aspect (i.e., a varying number of segments across INN networks, $M$) of INN needs to be investigated.
  • Convergence of INN for training can be proved mathematically.
  • Theoretical studies on the generalization of INN as a graph neural network (GNN) are needed.
  • INN training on ensembles of training data might be a way to enhance the generality of the model for image problems or noisy data.
We expect INN can open a next-generation Engineering Software 2.0 that not only merges separate modules of training, solving, and calibration but also tackles extremely challenging problems in the science and engineering domain.

### v2

@@
 Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration
+Engineering Software 2.0 by Interpolating Neural Networks: Unifying Training, Solving, and Calibration
+Chanwook Park
+Department of Mechanical Engineering
+Northwestern University
+Evanston, IL 60208
+chanwookpark2024@u.northwestern.edu
+&Sourav Saha
+Theoretical and Applied Mechanics Program
+Northwestern University
+Evanston, IL 60208
+sourav023@u.northwestern.edu
+&Jiachen Guo
+Theoretical and Applied Mechanics Program
+Northwestern University
+Evanston, IL 60208
+jiachenguo2020@u.northwestern.edu
+&Xiaoyu Xie
+Department of Mechanical Engineering
+Northwestern University
+Evanston, IL 60208
+xiaoyuxie2020@u.northwestern.edu
+&Satyajit Mojumder
+Theoretical and Applied Mechanics Program
+Northwestern University
+Evanston, IL 60208
+satyajitmojumder2022@u.northwestern.edu
+&Yangfan Li
+Department of Mechanical Engineering
+Northwestern University
+Evanston, IL 60208
+YangfanLi2024@u.northwestern.edu
+&Stefan Knapik
+Department of Mechanical Engineering
+Northwestern University
+Evanston, IL 60208
+stefan.knapik@u.northwestern.edu
+&Miguel A. Bessa
+School of Engineering
+Brown University
+Providence, RI
+miguel_bessa@brown.edu
+&Dong Qian
+Department of Mechanical Engineering
+University of Texas at Dallas
+Richardson, TX 75080
+dong.qian@utdallas.edu
+&Wei Chen
+Department of Mechanical Engineering
+Northwestern University
+Evanston, IL 60208
+weichen@northwestern.edu
+&Gregory J. Wagner
+Department of Mechanical Engineering
+Northwestern University
+Evanston, IL 60208
+gregory.wagner@northwestern.edu
+&Jian Cao
+Department of Mechanical Engineering
+Northwestern University
+Evanston, IL 60208
+jcao@northwestern.edu
+&Wing Kam Liu
+Department of Mechanical Engineering
+Northwestern University
+Evanston, IL 60208
+w-liu@northwestern.edu
+Equal contribution with Coauthor 1 and Coauthor 2Corresponding author
 Abstract
@@
 However, the application of these technologies within the field of engineering and science introduces distinct challenges.
-These challenges include scarcity and multi-modal aspect of data, low generalization error compared to classical numerical methods [4, 5, 6], slow inference (function evaluation) time that challenges real-time control [7, 8, 9, 10] and statistical modeling, and resource-heavy model architecture [11, 12, 13, 14].
+These challenges include scarcity and multi-modal aspect of data, low generalization error compared to classical numerical methods [4, 5, 6, 7], slow inference (function evaluation) time that challenges real-time control [8, 9, 10, 11] and statistical modeling, and resource-heavy model architecture [12, 13, 14, 15].
 In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of computational and data-driven engineering software into a new paradigm.
@@
 2 INN: Unifying Training/Learning, Solving, and Calibration
-Interpolating Neural Network (INN) is a systematically pruned neural network based on interpolation theories [17, 22] that leverage classical numerical methods such as finite element [23] and meshfree methods [24, 25].
+Interpolating Neural Network (INN) is a systematically pruned neural network based on interpolation theories [18, 7] that leverage classical numerical methods such as finite element [23, 24, 25] and meshfree methods [26, 27].
 Although these numerical methods and the neural network theory have been developed by seemingly different groups of researchers, they share one thing in common; the goal is to find a function.
@@
 The benchmark problems include training of data that follow scientific equations, classification of spiral data, classification of the fashion-MNIST dataset [35], calibration of parameters in the scientific equations and electrode manufacturing data [40], and solving of a parameterized partial differential equation (PDE).
-Except for the fashion-MNIST classification, all benchmark demonstrations are provided in the supplementary information, clearly revealing significant benefits such as less trainable parameters, high accuracy, less memory requirement, fast training, and inference time.
-One can also find how the INN hyperparameters are selected in Supplementary Information.
+Except for the fashion-MNIST classification, all benchmark demonstrations are provided in the supplementary information, clearly revealing benefits of INN such as less trainable parameters, high accuracy, less memory requirement, and fast training and inference time.
+One can also find how the INN hyperparameters are selected.
 Table 1 shows the benchmark result of the fashion-MNIST image classification problem [35].
@@
 Gamma:10 Max depth: 100 8$\times$8 pixels neurons 1 segments stride: 1
+Trainable N/A N/A N/A 535,818 470,400 288,000
+Parameters
 Test 0.8970 0.8730 0.8343 0.8918 0.8813 0.8731